the area of a certain state (state A) is seven times the area of another state (state B). If the sum of their areas is 8,000,000 sq miles, find the area of each state.
Answer by TimothyLamb(4379)

x = area of state A
y = area of state B
---
x = 7y
x + y = 8000000
---
put the system of linear equations into standard form
---
x - 7y = 0
x + y = 8000000
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = area of state A = 7000000 sq.mi.
y = area of state B = 1000000 sq.mi.
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---